OpenStudy (anonymous):

what is the equation of the normal to the curve: f(x)=xsquare 2 -5 at x=2

OpenStudy (anonymous):

\[f(x)=x^2-5\]?

OpenStudy (anonymous):

yes or no?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[f'(x)=2x\] Tangent line = y=4x-9 Normal line is ortangonal y=-1/4x+b -1=-1/2+b b=-1+1/2 b=-1/2 y=-1/4x -1/2